OpenStudy (anonymous):

Tiara wants to divide a square piece of paper into two equalivalent triangles. If the square measures 20 cm on each side, what will the third side of each triangle measure?

OpenStudy (anonymous):

A. 8.9 cm B. 20 cm C. 40 cm D. 28.3 cm

OpenStudy (anonymous):

 now use pythagorean theorem \[a^2+b^2=c^2\] \[20^2+20^2=d^2\]